OpenStudy (anonymous):

A certain quantity grows exponentially over time. The initial quantity at t = 0 is 100. The quantity grows by a factor of 2.5%. What is the quantity at t = 5? A. 0.0000009766 B. 88.1096 C. 113.1408 D. 305.1758

OpenStudy (anonymous):

What

OpenStudy (mokeira):

\[P+(1+r)^n\]

OpenStudy (mokeira):

this is what i meant

OpenStudy (anonymous):

yeah but how do i do it

OpenStudy (anonymous):

it grows by 2.5% (or *1.025) 5 times

OpenStudy (anonymous):

you mean r = 0.025

OpenStudy (anonymous):

is it C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

because 100*1.025*1.025*1.025*1.025*1.025 = 100*(1+0.025)^5 = 113.1408